A particle is moving in a one-dimensional box of width 10 Å. Calculate the probability of finding

the particle within an interval of 2 Å at the centre of the box when it is in the state of least energy


1
Expert's answer
2021-05-06T07:40:13-0400

Width of box, "l=10\space A^o"

Interval, "x=2\space A^o"

Since the electron is in its least energy state, "n=1"

Probability function, "\Psi_m=\sqrt{\dfrac{2}{l}}\space sin\bigg(\dfrac{n\pi x}{l}\bigg)"

"\Psi_m=\sqrt{\dfrac{2}{10}}\space sin\bigg(\dfrac{\pi\times 2}{10}\bigg)"

"\Psi_m=0.138"